# ORMNet: Object-centric Relationship Modeling for Egocentric Hand-object Segmentation

### v1

@@ III-F Training and inference
IV Experiments
This section first introduces the datasets and metrics we used to validate the EgoHOS performance. Second, we compare the proposed ORMNet with state-of-the-art methods in the EgoHOS task. Then, we ablate the import design strategy and elements of ORMNet.
IV-A Datasets and Metrics
EgoHOS. The EgoHOS [4] dataset consists of 11,743 egocentric images containing per-pixel segmentation labels of hands and interacting objects, gathered from Ego4D[55], EPICKITCHEN[56], THU-READ[57], and their own collected egocentric videos. Among them, 8,993 images are used for training, 1,124 images are for validation, 1,126 images are used for in-domain testing, and 500 images are used for out-of-domain testing.
mini-HOI4D. HOI4D[58] is a large-scale egocentric image and video dataset with rich annotations for category-level hand-object interaction. Frame-wise annotations for action segmentation, motion segmentation, panoptic segmentation, 3D hand pose, and category-level object pose are also provided, together with reconstructed object meshes and scene point cloud. In order to test the generalization of the proposed ORMNet, this paper further tests the proposed model with other comparative methods on the HOI4D dataset. Since the HOI4D dataset includes the masks of the hand and all objects that have interacted with the hand in a video, we first extract the masks of the objects currently interacting with the hand and delete other objects’ masks. Then, we map the label of the interacting objects to the left-hand objects, right-hand objects, and two-hand objects. Finally, we sampled a total of 1,095 egocentric images, forming a new dataset mini-HOI4D, to test the generalization of the algorithms on the EgoHOS task.
Following most related works, this paper uses the most common and widely used metrics in the segmentation task to evaluate the performance, i.e., we use Intersection-over-Union (IoU) and pixel Accuracy (Acc) for each category. Mean IoU (mIoU) and mean accuracy (mAcc) are also employed to indicate the overall performance.
IV-B Implementation Details
The experiments were conducted on 4 NVIDIA RTX 6000 Ada GPUs. The batch size was set to 12. Data preparation included cropping the image to (448, 448) and normalization with the mean of [106.011, 95.400, 87.429], and the standard deviation was set to [64.357, 60.889, 61.419]. The patch size of the Swin transformer encoder was set to 4, and the window size was set to 12. The AdamW optimizer was used in the experiment, the learning rate was set to 1e-5, and the weight decay was set to 0.01.
IV-C Quantitative Results
@@
1) Comparison results on EgoHOS in-domain test set. Table I presents the comparison results of our proposed ORMNet against state-of-the-art methods on the EgoHOS in-domain test set. In addition to the parallel (Para.-series) and sequential methods (Seq.-series) proposed in Zhang’s work[4] introduced in detail in Sec. II-B, we also select some of the latest transformer-based models with different backbones as parallel methods to compare the effectiveness of the model, e.g., Segformer[13], Segmenter[37], MaskFormer[38], and Mask2former[39], and SCTNet[54]. All these methods are trained on the EgoHOS train set and tested on the in-domain test set using the same parameter setting as the proposed ORMNet. In general, the proposed ORMNet far outperforms all previous methods in most categories on both metrics. Specifically, compared with the previous best parallel method in the $8^{th}$ row, Mask2Former, our method has demonstrated pronounced advancements across various categories and exhibited particularly notable enhancements in object segmentation. For example, our method showcases substantial enhancements of 15.85%, 10.64%, and 3.6% on left-, right-, and two-hand objects, respectively. Furthermore, compared with the results of the previous best sequential methods $\text{Seq}^{\flat}$, our ORMNet improves IoU by more than 4% on both the left and right hands. The most substantial enhancement is observed in the right-hand objects, surpassing 12%. Finally, compared with the previous best method, our proposed method attains a commendable advancement of 4.33% and 4.55% in mIoU and mAcc, respectively. This is because we propose an ORD strategy to decouple the relationship between object categories, which reduces the burden of network learning. Furthermore, we propose the HOR module to strengthen the correlation between hand features and object features, significantly improving the segmentation accuracy of hands and objects.
TABLE II: Comparison results on the egohos out-of-domain test set.
@@
TABLE III: Comparison results on the mini-HOI4D test set. Note: The left-hand objects are not considered to compare due to few samples.
@@
2) Comparison results on EgoHOS out-of-domain test set. In order to test the generalization performance of different models, the EgoHOS dataset provides 500 out-of-domain images, which are captured from YouTube and annotated manually. Table II illustrates the comparison results of our method with the previous methods on the EgoHOS out-of-domain test set measured by IoU and Acc. Please note that some of the results in the table are from direct inference using the officially released best model (if available), and other results are from direct inference on the best model trained on our own. It can be observed that the results of our proposed ORMNet on the two metrics of most categories have obtained competitive generalization performance. Among them, the IoU of the left and right hand exceed 94%, and the Acc. of these two categories exceed 96%. The segmentation results on various object categories also achieve balanced performance. Finally, our ORMNet achieves state-of-the-art performance in overall mIoU and mAcc metrics and demonstrates excellent generalization performance on the out-of-domain test set.
3) Comparison results on mini-HOI4D dataset. In addition to the evaluation on the EgoHOS out-of-domain test set, we also assess the model’s generalization capabilities using a completely different mini-HOI4D dataset. It should be noted that due to the limited samples of left-hand objects in the original HOI4D[58] dataset, the generated mini-HOI4D dataset does not include samples of left-hand objects. Consequently, no testing is performed on left-hand objects.
The comparison results on the mini-HOI4D data set are shown in Table III. Similarly to the EgoHOS out-of-domain test set, the results on the mini-HOI4D dataset are directly inferred by utilizing trained and saved checkpoints without further training procedures. Overall, our proposed ORMNet outperforms all the previous parallel and sequential methods concerning mIoU and mAcc. Our method excels notably in hand segmentation, i.e., 1.57% and 3.71% improvements on IoU and Acc for the right hand compared with the $8^{th}$ row. Our method also competes well in interacting object segmentation, yielding commendable outcomes. Therefore, it can be deduced from the results of the EgoHOS out-of-domain and mini-HOI4D test sets that our proposed ORMNet exhibits superior generalization capabilities in contrast to the comparative methods.
[…]
TABLE IV: Ablation results on the EgoHos in-domain test set under the same setting measured by IoU and mIoU. Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
@@
TABLE V: The experimental results using different hyper-parameters. $\alpha$, $\gamma$, and $\lambda$ are the weights of objects, hands, and CB losses. Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
@@
4) Ablation study. In order to verify the effectiveness of the proposed method alongside its corresponding sub-module and strategy, we conduct the ablation study in this section. The experimental results are shown in Table IV. The basic model uses Swin-B as the encoder and uses three decoder branches to predict hands, contact boundaries, and objects. This basic model is a naive solution for the EgoHOS task and does not model any relationships between hands and interacting objects, so its results are relatively poor. The $2^{nd}$ and $3^{rd}$ rows of the table show the results after adding the ORD strategy and HOR module independently, which proves that adding them to the basic model is beneficial. The last row is the result after integrating the ORD strategy and the HOR module. By explicitly modeling both the hand-object relationship and the object-object relationship, the segmentation accuracy achieves its peak performance. Generally speaking, although the effect improvement brought by the ORD strategy is higher than that of the HOR module, using these two modules and strategies simultaneously can further improve the segmentation performance, which verifies the effectiveness of our proposed method and sub-methods.
Figure 5: Visualization results of the ORMNet compared with the sequential method on the EgoHOS out-of-domain test set (left) and mini-HOI4D dataset (right). Red: left hand, dark blue: right hand, pink: left-hand objects, light blue: right-hand objects, and green: two-hand objects.
5) Hyper-parameter experiments. While identifying the optimal hyper-parameters is not the primary focus of this paper, this section still conducts experiments on the configurations of different hyper-parameters. The experimental results are shown in Table V. We assign the value of $\alpha$, $\gamma$, and $\lambda$ based on the relative significance of the interacting objects, hands, and contact boundaries. In the $1^{st}$ row, all categories are considered equal importance, while in subsequent experiments (i.e., $2^{nd},3^{rd}$ and $4^{th}$), we regard the objects, CBs, and hands as the most crucial, respectively. The results lead to the inference that amplifying the importance of hands can yield enhancements in the method’s performance, which can be explained intuitively because the proposed HOR module leverages hands as prior knowledge for feature enhancement.
IV-D Qualitative Results
@@
2) Results on EgoHOS out-of-domain test set. We directly used the saved checkpoint to conduct experiments on the EgoHOS out-of-domain test set and compared it with the $\text{Seq}^{\flat}$ method. The visualized experimental results are shown in Fig. 5 (left). According to the visualization results, our ORMNet has stronger generalization than the compared sequential model. Specifically, our method can segment and identify left and right hands accurately, as shown in the $3^{rd}$ to the last rows of results. Furthermore, our method can precisely segment and recognize different kinds of objects, ensuring consistent segmentation of interactive objects.
3) Results on mini-HOI4D dataset. The experiments on the mini-HOI4D dataset are also conducted directly using the saved checkpoints. The experimental results are shown in Fig. 5 (right). It can be observed from the experimental results that the contrasting sequential method easily confuses the left and right hands, such as $1^{st}$, $2^{nd}$, $4^{th}$, and $5^{th}$ rows of the results. In contrast, our method exhibits robust generalization capabilities coupled with outstanding segmentation performance. In addition, although some tiny parts are not considered, such as the handle of the bucket in the third row, our method also significantly excels in segmenting two-hand objects.
V Conclusion
This paper introduces an end-to-end high-performance Object-centric Relationship Modeling Network to effectively handle the egocentric hand-object segmentation task. First, the HOR module is proposed to facilitate the modeling of hand-object interactions through hand-guided cross-attention, enhancing the network’s adaptability to learn precise representations. Second, this paper designs the ORD strategy to decouple the two-hand objects, thereby reducing the network’s confusion about object segmentation. Experimental results demonstrate that the proposed ORMNet exhibits SOTA segmentation efficacy and robust generalization capabilities across varied datasets.
VI Acknowledgements

### v2

@@ 4.3.1 In-domain Results
Comparison results on EgoHOS in-domain test set. To validate the effectiveness of the proposed ORMNet, we conducted experiments by training on the EgoHOS train set and testing on the EgoHOS in-domain test set. Due to the novelty of the EgoHOS task, we selected a set of novel state-of-the-art general segmentation algorithms [16, 46, 48, 49, 45] as compared parallel methods. Furthermore, we ensured a comprehensive comparison by incorporating all existing methods [9] tailored for the EgoHOS task in this evaluation. The results are presented in Table 1. In general, the proposed ORMNet far outperforms all previous methods in most categories on both metrics. Specifically, compared with the previous best parallel method - Mask2Former [49], our method has demonstrated pronounced advancements across various categories and exhibited particularly notable enhancements in interacting-object segmentation. For example, our method showcases substantial enhancements of 15.85%, 10.64%, and 3.6% on left-, right-, and two-hand objects, respectively. Moreover, compared with the results of the previous best sequential methods $\text{Seq}^{\flat}$, our ORMNet improves IoU by more than 4% on both the left and right hands. The most substantial enhancement is observed in the right-hand objects, surpassing 12%. Finally, compared with the previous best method, our proposed method attains a commendable advancement of 4.33% and 4.55% in mIoU and mAcc, respectively. These results profoundly demonstrate the state-of-the-art segmentation performance achieved by the proposed ORMNet on the basis of end-to-end optimization, leveraging the HOR module and the ORD strategy to model the correlations between hands and objects as well as objects and objects.
Table 2: Comparison results on the EgoHOS out-of-domain test set. The blue letters represent the best results. The ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
@@
4.3.2 Out-of-domain Results
1) Comparison results on EgoHOS out-of-domain test set. In order to test the generalization performance of different models, the EgoHOS dataset provides 500 out-of-domain images, which are captured from YouTube and annotated manually. Thus, We compare ORMNet with different parallel and sequential models by directly testing saved best models on the EgoHOS out-of-domain test set, results are shown in Table 2. It can be observed that the results of our proposed ORMNet on the two metrics of most categories have obtained significant generalization performance. Among them, the IoU of the left and right hand exceed 94%, and the Acc. of these two categories exceed 96%. The segmentation results on various interacting-object categories also achieve balanced performance. Finally, our ORMNet achieves state-of-the-art generalization performance in overall mIoU and mAcc metrics on the EgoHOS out-of-domain test set.
Table 3: Comparison results on the mini-HOI4D test set. The blue letters represent the best results. The ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
@@ 4.3.2 Out-of-domain Results
2) Comparison results on mini-HOI4D dataset. In order to comprehensively validate the generalizability of the proposed method, we also assess the generalization capabilities of algorithms using a completely different mini-HOI4D dataset. It should be noted that since the original HOI4D [43] dataset contains very few samples of left-hand objects, we do not perform testing on left-hand object category. We also utilized the saved best models to test on the mini-HOI4D dataset without further training procedures, the comparison results are shown in Table 3. Overall, our proposed ORMNet outperforms all the previous parallel and sequential methods concerning mIoU and mAcc. Our method excels notably in hand segmentation, i.e., 1.57% and 3.71% improvements on IoU and Acc for the right hand compared with the previous best Mask2Former model. Our method also competes well in interacting-object segmentation, yielding commendable outcomes. Therefore, it can be deduced from the results of the EgoHOS out-of-domain and mini-HOI4D test sets that our proposed ORMNet exhibits superior generalization capabilities in contrast to the comparative methods.
Table 4: Ablation results on the EgoHOS in-domain test set under the same setting measured by IoU and mIoU. Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
@@
4.3.3 Ablation Study
In order to verify the effectiveness of the proposed method alongside its corresponding sub-module and strategy, we conduct the ablation study in this section. The experimental results are shown in Table 4. The basic model uses Swin-B as the encoder and uses three decoder branches to predict hands, contact boundaries, and interacting objects. This basic model is a naive solution for the EgoHOS task and does not model any relationships between hands and interacting objects, so its results are relatively poor. The $2^{nd}$ and $3^{rd}$ rows of the table show the results after adding the ORD strategy and HOR module independently, which proves that adding them to the basic model is beneficial. The last row is the result after integrating the ORD strategy and the HOR module. By explicitly modeling the relations between hands and objects as well as objects and objects, the segmentation accuracy achieves its peak performance. Generally speaking, although the effect improvement brought by the ORD strategy is higher than that of the HOR module, using these two modules and strategies simultaneously can further improve the segmentation performance, which verifies the effectiveness of our proposed method and sub-methods.
4.3.4 Hyper-parameter Study
The method presented in this paper requires setting multiple hyper-parameters for total training loss, i.e., $\alpha,\gamma$, and $\lambda$. This section conducts experiments on the configurations of different hyper-parameters, results are shown in Table 5. We assign the values of $\alpha$, $\gamma$, and $\lambda$ based on the relative significance of the interacting objects, hands, and contact boundaries. In the $1^{st}$ row, all categories are considered equal importance, while in subsequent experiments (i.e., $2^{nd},3^{rd}$ and $4^{th}$), we regard the interacting objects, CBs, and hands as the most crucial, respectively. The results lead to the inference that amplifying the importance of hands can yield enhancements in the method’s performance, which can be explained intuitively because the proposed HOR module leverages hands as prior knowledge for interacting-object feature enhancement.
Table 5: The experimental results using different hyper-parameters. $\alpha$, $\gamma$, and $\lambda$ are the weights of interacting objects, hands, and CB losses.
@@ 4.3.4 Hyper-parameter Study
4.4 Qualitative Results
4.4.1 In-domain Results
@@
Figure 5: Visualization results of two-hand objects on the EgoHOS in-domain test set compared with sequential method $\text{Seq}^{\flat}$ [9].
[…]
Figure 6: Visualization results of the ORMNet compared with the sequential method [9] on the EgoHOS out-of-domain test set (left) and mini-HOI4D dataset (right). The main improvements are highlighted in the dashed yellow box.
4.4.2 Out-of-domain Results
1) Results on EgoHOS out-of-domain test set. We visualize the comparative results on the EgoHOS out-of-domain test set with the $\text{Seq}^{\flat}$ method in Fig. 6 (left). The figure shows that our ORMNet has stronger generalization than the compared sequential model. Specifically, our method can segment and identify left and right hands accurately, as shown in the $3^{rd}$ to the last rows of results. Furthermore, our method can precisely segment and recognize different kinds of interacting objects, as shown in the first and second rows in the figure.
2) Results on mini-HOI4D dataset. We visualize the comparative results on the mini-HOI4D test set with the $\text{Seq}^{\flat}$ method in Fig. 6 (right). It can be observed that the contrasting sequential method easily confuses the left and right hands, such as $1^{st}$, $2^{nd}$, $4^{th}$, and $5^{th}$ rows of the results. In contrast, our method exhibits robust generalization capabilities coupled with outstanding segmentation performance.
5 Conclusion
This paper introduces the ORMNet to effectively handle the EgoHOS task while achieving end-to-end optimization by incorporating two parts. First, the HOR module is proposed to facilitate the modeling of hand-object interactions through hand-guided cross-attention, enhancing the network’s adaptability to learn comprehensive interacting-object representations. Second, this paper designs an ingenious ORD strategy to decouple the two-hand objects, thereby reducing the confusion of the network about interacting-object classification and emphasizing the left/right interactions.
However, there remains a limitation that needs to be addressed. Specifically, the EgoHOS task focuses solely on predicting hands (left- and right-hand) and interacting objects (left-hand object, right-hand object, and two-hand object) without inferring the object’s name or semantic information. Consequently, this limits the comprehensive description of people’s actions. To address this issue, we plan to enhance and refine the EgoHOS task in future work. Our goal is to incorporate the prediction of interacting-object semantic information alongside hand-object segmentation, thereby expanding the practical applicability of the algorithm in real-world scenarios.
6 Acknowledgements